# Beyond Labeling Oracles: What does it mean to steal ML models?

### v1

@@ Abstract
Model extraction attacks are designed to steal trained models with only query access, as is often provided through APIs that ML-as-a-Service providers offer. ML models are expensive to train, in part because data is hard to obtain, and a primary incentive for model extraction is to acquire a model while incurring less cost than training from scratch. Literature on model extraction commonly claims or presumes that the attacker is able to save on both data acquisition and labeling costs. We show that the attacker often does not. This is because current attacks implicitly rely on the adversary being able to sample from the victim model’s data distribution. We thoroughly evaluate factors influencing the success of model extraction. We discover that prior knowledge of the attacker, i.e. access to in-distribution data, dominates other factors like the attack policy the adversary follows to choose which queries to make to the victim model API. Thus, an adversary looking to develop an equally capable model with a fixed budget has little practical incentive to perform model extraction, since for the attack to work they need to collect in-distribution data, saving only on the cost of labeling. With low labeling costs in the current market, the usefulness of such attacks is questionable. Ultimately, we demonstrate that the effect of prior knowledge needs to be explicitly decoupled from the attack policy. To this end, we propose a benchmark to evaluate attack policy directly.
@@ 1 Introduction
Modern ML models are valuable intellectual property, in part because they are expensive to train (Sharir et al., 2020; Patterson et al., 2021), and model extraction (ME) attacks (also known as model stealing or black-box estimation) threaten their confidentiality. In a ME attack, the adversary attempts to steal a victim model over query access, in order to obtain an approximate copy of that model with similar performance (Tramèr et al., 2016). In the most common class of ME attacks, the adversary uses queries to the victim model as a training set for the copy of that model. Figure 5 illustrates such a typical attack, while we discuss other types of ME attacks in Section 2.1.
@@
ME attacks are a looming threat for ML-as-a-Service providers, whose business model depends on paid query access to their models over APIs. In fact, Kumar et al. (2020) outline that ME is one of the most concerning threats for the industry. Typically, the claimed motivation to conduct ME attacks is to avoid the costs (see Appendix A) involved in training a model from scratch; the implicit assumption is that ME is data- and/or compute- cheaper (Jagielski et al., 2020). Here, the adversary’s goal is to get the best performing model using the least resources. Current literature uses the ratio between the number of samples $M$ used to train the victim model and the number of queries $N$ used to train the stolen copy as a measurement of the attack’s efficacy; If $M$ is much larger than $N$, the attack is a success. This criterion stems from the costs associated with data labeling and training.
Intuitively, this query efficiency criterion can be possible, provided the attacker can identify the most informative queries.
However, we find that much of current research does not adequately account for how the adversary’s prior knowledge contributes to the attacks’ query efficiency. In many experiments, prior work assumes that the attacker has access to unlabeled data from a similar distribution to the victim’s training data, but does not rigorously study how having this prior to the victim’s training distribution contributes to the attack’s efficiency (Correia-Silva et al., 2018; Orekondy et al., 2019a; Dziedzic et al., 2022a; Jagielski et al., 2020; Pal et al., 2020; Zhang et al., 2021; Okada et al., 2020; Karmakar and Basu, 2023). We refer to these as data access assumptions. In contrast, when such priors are removed, the query complexity of data-free model extraction attacks drastically increases (Truong et al., 2021). This is important because, given enough prior knowledge of in-distribution data, it becomes cheaper for the adversary to train their own model from scratch without querying the victim. If all the adversary needs is labels for the unlabeled data they have access to, in many use cases it might be more cost-effective to obtain those using a crowdsourcing service, e.g. Amazon’s Mechanical Turk, than by querying the victim model – as we discuss in Appendix A. This, in turn, removes their incentive to perform ME.
In this paper, we study how data access assumptions influence the accuracy of extracted models, and how this changes adversaries’ incentives. First, we show that with the amount of prior knowledge (i.e. unlabeled datapoints) that many current works assume, for commonly used benchmarks one can train a model with comparable accuracy to that of the victim model. Second, we demonstrate that soft-labels only lead to minor improvements in accuracy of the stolen model ($\sim 4\%$) for cases with little prior knowledge, and the effect disappears with more data available. In other words, aside from providing a label, victim labels leak limited information about the decision boundaries.
Next, to better understand the adversary’s dependence on prior knowledge of the distribution, we investigate a surprising phenomenon where models can sometimes get extracted with data that is unrelated to the model task (Orekondy et al., 2019a). We investigate the extent to which such out-of-distribution (OOD) queries expose the model’s behaviour for in-distribution (IND) data. We suggest a novel benchmark with explicit control over the informativeness of OOD queries by instructing the model to give fake, yet plausible, decisions. We show that this causes both a drop of the stolen model accuracy and increases data complexity. Furthermore, by reducing informativeness of OOD data, we implicitly force an intelligent adversary to solve a problem of predicting if a given datapoint is IND or OOD, a problem known to be as hard as classification itself (Tramer, 2021). Put simply, to extract a model in this setting, an attacker needs to either have IND data, or a very large query budget. Yet, when such data is available, the attacker once more gets nothing but a label from the victim.
Overall, our paper finds that the practical threat posed by ME attacks is often exaggerated, since the attacker bears comparable costs to the victim. To summarize, we make the following contributions:
@@
  • We examine how prior knowledge of the victim model’s training distribution contributes to the efficiency of ME. We show that if the adversary has access to even just a small amount of IND data, ME simply becomes a labeling oracle. In other words, aside from providing a label, victim labels leak limited information about the decision boundaries.
  • We then investigate cases where OOD data is used for ME and find that IND data dominates extraction, yet OOD also leaks information about IND decision boundaries (Section 5).
  • We modify victim models so that, for OOD queries, they output labels that are uncorrelated with IND decision boundaries, disrupting ME (Section 5). Our new benchmark controls for IND leakage from OOD queries and evaluates the attacker’s knowledge of the distribution.
@@ 2 Related work
2.1 Model Extraction Attacks
Model extraction attacks, also called model stealing, were first explored by Tramèr et al., who proposed an attacker that queries the victim model to label its dataset and trains a model to match these predictions. Various works extended the attack with the goal of achieving similar task accuracy while minimizing the number of queries required. Most of these attacks either assume access to a surrogate dataset (Correia-Silva et al., 2018; Orekondy et al., 2019a) or to a portion of the real training set (Rakin et al., 2021), or use random queries (Krishna et al., 2019; Chandrasekaran et al., 2020) in domains like NLP. We discuss the relation between ME to active learning in Appendix B. Truong et al. attested that, in order to successfully extract the victim model, the attacker’s dataset must share semantic or distributional similarity to the real training dataset; otherwise resulting in an insufficient extraction accuracy. To mitigate this issue, they propose a data-free model extraction attack (DFME) that does not require a surrogate dataset. Inspired by data-free knowledge distillation (Lopes et al., 2017), they train a generative model to synthesize queries which maximize disagreement between the attacker’s and the victim’s models. A similar method was also proposed by MAZE (Kariyappa et al., 2021). These results align with our analysis and show that the adversary can compensate for the absence of prior knowledge by using a very large query budget.
Jagielski et al. proposed a ME attack that, in addition to accuracy, targets fidelity. In this scenario, the attacker focuses on the exact reproduction of the victim model behaviour on all possible inputs. Carlini et al. argued that ME attacks are essentially a cryptanalytic problem, and proposed an attack based on differential cryptanalysis, with the goal of high fidelity extraction. Note that this line of research falls outside the scope of our work, since both the worst and the average-case query complexities reported by both papers significantly exceed accuracy-based ME attacks empirically. As such, in this work we will focus our evaluations on ME attacks that target task accuracy.
@@ 2.2 Defenses Against Model Extraction Attacks
Having established the vulnerability of ML models to ME attacks, we turn to defenses. It is worth mentioning that in the limit there is little that a victim can do to stop ME. Unlimited data sampling allows for exhaustive search of all possible inputs, whereas functions that are used to approximate decision boundaries in the limit can approximate arbitrary functions (Hornik et al., 1989). This leads to an inherent performance trade-off, where the victim sacrifices performance of the model to limit information leakage to an acceptable level. Broadly, defenses against ME attacks can be categorized as active, passive, reactive, or proactive, based on when they are applied in the extraction process.
Active defenses are applied to the model’s predictions with the purpose of manipulating the prediction of the attacker’s queries. This includes prediction poisoning (Orekondy et al., 2019b), where small perturbations are added to the model’s predictions to poison the training objective of the attacker, and prediction truncation (Tramèr et al., 2016) These defenses come at the cost of lowering the quality of predictions for legitimate users. Passive defenses attempt to detect and block an attack (Juuti et al., 2019). Reactive defenses address extraction attacks post-hoc by determining if a suspect model was stolen or not. This includes watermarking (Jia et al., 2021a), dataset inference (Maini et al., 2021), and proof-of-learning (Jia et al., 2021b). Dziedzic et al. proposed a proactive defense, which attempts to deter the attack before it happens, by requiring the completion of a proof-of-work.
We argue that, although our OOD component degrades ME performance, it should not be considered a defense mechanism. Rather, we position it as a benchmark for validating our hypothesis regarding the adversary’s assumption over the informativeness of the OOD region. It does, however, share a common intuition with many passive and active defenses, in the sense that they too target the amount of information that can be obtained from querying the victim model, across the entire input space.
3 Warm-up: Linear classification
@@
In this section, we elaborate on our intuition behind the success of model extraction attacks, which later motivates the design of our benchmark to evaluate the impact of prior knowledge on ME attacks.
@@
Consider a linear classification model that splits the input space into two decision regions, and the corresponding decision boundary is defined by the line $y=\alpha x$. Here, $(x,y)$ denotes the two-dimensional input, and $\alpha$ is the single model parameter that determines the entire decision boundary, which we demonstrate in Figure 1. ME in this context is equivalent to estimating the decision boundary over the green IND region where $x\sim[20,80]$. The attacker is assumed to have the ability to query the model with arbitrary data but has no knowledge of where the IND region starts or ends. We assume that the depicted red region, $x\sim[0,20]\cup[80,100]$, is OOD. Note that it is only important to perform well on the task of interest in the IND region, and neither the victim model’s owner nor the attacker are concerned with making predictions for OOD inputs. An attacker with no precise knowledge of the range of the input domain, i.e. what is IND and what is out of it, will require more parameters to approximate the whole range, since they do not know where the task begins and ends.
@@
The left-most plot represents the case of a naive victim model with the original decision boundary. Here ME reduces to estimating a single parameter $\alpha$. Note that, in this case, the IND model can be extracted equally well using both IND or OOD data points, as the model behaviour is shared between both domains. In the middle case, we slightly increase the complexity of the decision boundary in the OOD area by introducing a new linear boundary with non-zero intercept in each region. As the attacker does not know where the IND and OOD regions are, it must approximate a piece-wise linear boundary, i.e. the slopes and the intercepts of the new lines, in addition to the slope of the original IND boundary. This increases the required number of parameters to 5 and, consequently, the cost of ME. In the right-most case, we further increase the complexity as before by introducing two additional linear decision boundaries per region, resulting in a total of 9 parameters to be estimated. The above example demonstrates that while the IND area was not modified, adding complex and task-independent boundaries to OOD regions can significantly increase the attacker’s required capacity, thus making model extraction more costly. The same intuition also holds for the minimal required sample size. In the left-most case, the adversary needs at least 2 data points to estimate the parameter $\alpha$; in the middle case, the adversary needs at least 6 data points; and for the right-most case, at least 10 data points are required. The increase in the complexity of the decision boundary to be estimated results in an increase in both capacity and the minimal required sample size.
@@
This behaviour is related to the locally independent nature of large models when performing regression or classification tasks. As sampling the boundary at one point does not necessarily reveal any information on the boundary IND, the attacker must explore the entire space. The main assumption here is that the attacker can not know which areas are important. In other words, we find that model extraction attacks implicitly assume that the adversary has prior knowledge of the distribution to be able to reproduce the victim model’s predictions on the task of interest. Without such knowledge, the adversary is unable to tell whether making a query to the victim model will aid its goal of learning the true decision boundaries. In Section 5.2, we evaluate this assumption by equipping the victim model with an OOD detection component. In this setting, we find that the adversaries end up “wasting” queries on learning non-important boundaries that do not contribute to their goal.
Hence, we demonstrate that ME adversaries either need access to prior knowledge about the distribution or the ability to submit a large number of queries to the victim model.
4 Model Extraction Attacks and Prior Knowledge
Refer to caption Refer to caption (a) CIFAR-10 Refer to caption (b) MNLI
Figure 2: Evaluation of the risk posed by an attacker with some prior knowledge over the true data distribution. The prior knowledge is expressed as access to a percentage of the true training set. Attackers with more than $50\%$ access can fully extract the model. However, in this case, the extraction is not more efficient than training from scratch. Attackers do not gain much by querying the victim model instead of using real labels, which also appears equivalent to label-only access to the victim. This shows that the victim is essentially a labeling oracle.
To evaluate the risk ME attacks pose, we measure the effect of the adversary’s prior knowledge over the data distribution. Section 4.2 presents a baseline attacker that only utilizes its prior knowledge of the data distribution to train a model. Section 4.3 evaluates commonly used ME attacks and demonstrates that they only marginally improve upon the baseline attack accuracy.
4.1 Experimental Setting

### v3

@@ Abstract
Model extraction attacks are designed to steal trained models with only query access, as is often provided through APIs that ML-as-a-Service providers offer. Machine Learning (ML) models are expensive to train, in part because data is hard to obtain, and a primary incentive for model extraction is to acquire a model while incurring less cost than training from scratch. Literature on model extraction commonly claims or presumes that the attacker is able to save on both data acquisition and labeling costs. We thoroughly evaluate this assumption and find that the attacker often does not. This is because current attacks implicitly rely on the adversary being able to sample from the victim model’s data distribution. We thoroughly research factors influencing the success of model extraction. We discover that prior knowledge of the attacker, i.e. access to in-distribution data, dominates other factors like the attack policy the adversary follows to choose which queries to make to the victim model API. Our findings urge the community to redefine the adversarial goals of ME attacks as current evaluation methods misinterpret the ME performance.
@@ 1 Introduction
Modern ML models are valuable intellectual property, in part because they are expensive to train (Sharir et al., 2020; Patterson et al., 2021), and model extraction (ME) attacks threaten their confidentiality. In a ME attack, the adversary attempts to steal a victim model over query access, in order to obtain an approximate copy of that model with similar performance (Tramèr et al., 2016). In the most common class of ME attacks, the adversary uses queries to the victim model as a training set for obtaining a copy of that model [^1] Figure 6 illustrates such a typical attack, while other types of ME attacks are discussed in Section 2..
@@
ME attacks are a looming threat for ML-as-a-Service providers, whose business model depends on paid query access to their models over APIs. In fact, Kumar et al. (2020) outlines that ME is one of the most concerning threats to the industry. Typically, the claimed motivation to conduct ME attacks is to avoid the costs involved in training a model from scratch. Most notably, this refers to the data collection and data labeling costs, and sometimes mentions the training computational costs. The implicit assumption is that ME is data- and/or compute- cheaper (Jagielski et al., 2020) – the adversary’s goal is to get the best performing model using the least resources. This includes using easily available queries, possibly from another distribution than the victim’s as well as minimizing the query complexity, i.e. stealing the model using less samples than was used to train it in the first place. Intuitively, this query efficiency criterion can be possible, provided the attacker identifies the most informative queries from the available query distribution.
In this paper we investigate primary assumptions and arguments made in the ME literature. We find that current research does not adequately account for how the adversary’s prior knowledge contributes to their ability to choose the right queries to make. Many existing attacks demonstrating impressive performance assume the attacker has access to either a small number of unlabeled samples from the true victim’s training distribution (thereafter, in-distribution data or IND), or a potentially larger number of unlabeled samples from a different, yet similar, distribution (thereafter, out-of-distribution data or OOD) (Correia-Silva et al., 2018; Orekondy et al., 2019a; Dziedzic et al., 2022; Jagielski et al., 2020; Pal et al., 2020; Zhang et al., 2021; Okada et al., 2020; Karmakar and Basu, 2023). In contrast, attacks that do not assume any prior knowledge, and perform in a data-prior free manner, usually exhibit a drastically higher query complexity (Truong et al., 2021; Jagielski et al., 2020).
We extensively experiment with ME attacks using additional control mechanisms to disambiguate the effect of prior knowledge from the attack policy. We find that performance of current ME attacks is dominated by the IND data; in settings where IND is mixed with OOD data to decrease the data costs, the impact of OOD is minor; moreover, when only IND is used, we find that the victim model merely serves as a labeling oracle, and not necessarily the most cost-effective one. Overall, our position is that the practical threat posed by ME attacks is often exaggerated since the attacker bears comparable costs to the victim.
To summarize, we make the following contributions:
  • We demonstrate how prior knowledge dominates ME performance and cost efficiency, even in cases where OOD data is used to augment the attacker’s query distribution, thus leaking information about IND decision boundaries at lower acquisition costs.
  • By modifying victim models such that the OOD leakage is reduced, we further demonstrate the dependence on prior knowledge, which can only be compensated by increasing the query complexity.
@@
  • We show that if the adversary has access to even just a small amount of IND data, ME becomes a labeling oracle. In other words, aside from providing a label, victim labels leak limited information about the decision boundaries.
@@ 2 Related work
Model extraction attacks, also called model stealing, were first explored by Tramèr et al., who proposed an attacker that queries the victim model to label its dataset and trains a model to match these predictions. Various works extended the attack with the goal of achieving similar task accuracy while minimizing the number of queries required. Most of these attacks either assume access to a surrogate dataset (Correia-Silva et al., 2018; Orekondy et al., 2019a) or to a portion of the real training set (Rakin et al., 2021), or use random queries (Krishna et al., 2019; Chandrasekaran et al., 2020) in domains like NLP. We discuss the relation between ME to active learning in Appendix A. Truong et al. attested that, in order to successfully extract the victim model, the attacker’s dataset must share semantic or distributional similarity to the real training dataset; otherwise resulting in an insufficient extraction accuracy. To mitigate this issue, they propose a data-free model extraction attack (DFME) that does not require a surrogate dataset. Inspired by data-free knowledge distillation (Lopes et al., 2017), they train a generative model to synthesize queries which maximize disagreement between the attacker’s and the victim’s models. A similar method was also proposed by MAZE (Kariyappa et al., 2021). (Lin et al., 2023) investigated ways to reduce the query complexity of data-free approaches. These results align with our analysis and show that the adversary can compensate for the absence of prior knowledge by using a very large query budget.
Jagielski et al. proposed a ME attack that, in addition to accuracy, targets fidelity. In this scenario, the attacker focuses on the exact reproduction of the victim model behaviour on all possible inputs up to symmetries. Carlini et al. argued that ME attacks are essentially a cryptanalytic problem, and proposed an attack based on differential cryptanalysis, with the goal of high fidelity extraction. Note that this line of research falls outside the scope of our work, since both the worst and the average-case query complexities reported by both papers significantly exceed accuracy-based ME attacks empirically. As such, in this work we will focus our evaluations on ME attacks that target task accuracy.
@@
It is worth mentioning that in the limit there is little that a victim can do to stop ME. Unlimited data sampling allows for exhaustive search of all possible inputs, whereas functions that are used to approximate decision boundaries in the limit can approximate arbitrary functions (Hornik et al., 1989). This leads to an inherent performance trade-off, where the victim sacrifices performance of the model to limit information leakage to an acceptable level.
3 Background
In this section, we first discuss the different costs involved in model extraction and connect them with the ME literature; we cover related works in detail in Section 2. We then describe how prior literature approached cost reduction using random data, why this can work theoretically, and what assumptions are required. Subsequently, we describe the additional mechanism we use to disambiguate the effect of the attack policy from increased query budget in ME.
3.1 Definitions
In what follows, we refer to in-distribution data as IND and out-of-distribution data as OOD. Note that we follow the conventions of the field and use these terms loosely to describe the distribution that training data comes from for the victim model as IND, while OOD refers to any data that comes from a different distribution.
We define an ME adversary as 3-tuple $(\mathcal{D}_{\sf IND},\mathcal{D}_{\sf OOD},\Pi)$, where $\mathcal{D}_{\sf IND}$ represents the adversary’s access to the IND, i.e. it’s prior knowledge, $\mathcal{D}_{\sf OOD}$ represents the adversary’s unrelated OOD query distribution, and $\Pi$ represents the adversary’s query selection policy. When attacking, the adversary uses $\Pi$ to select samples to be queried from its IND and OOD distributions. Then, the queries are issued to the victim model to obtain labels. These labels can be either full probability vectors, top-k probabilities, or label-only. In our evaluations, we focus on the stronger setting of full probability vectors. At last, the adversary trains a model on the labeled query set. Importantly, since attacks in the current literature vary all three from the above, it is hard to pinpoint why a given attack seemingly performs better – could it be a better attack policy or perhaps more informative data? This is the question we answer.
3.2 Primer on ML costing
We are not aware of any ME literature that formally defines what makes model extraction attacks successful. While high-fidelity extraction attackers explicitly expect to extract a model within a given error, intuitively, accuracy-driven ME can be considered successful if stealing a model costs less than developing it. That is, as long as the accuracy of the stolen model matches the victim it is a success. The overall cost of producing and deploying a model can be broken down into three main parts: (1) data collection, (2) data labeling, and (3) model training. We note that there are other costs corresponding to ML infrastructure, however they are not relevant to the current ME threat model.
Now we turn to CIFAR10 as a cost case study – note that CIFAR10 is one of the most evaluated benchmarks for ME. Here we assume CIFAR10 dataset with 60k samples that can be labeled for $35\$*50+10*25\$=2000\$$ with Google Cloud Cloud (2024) and $0.04\$*60000=2400\$$ with Amazon Sagemaker Amazon (2024). Data collection here is practically free, since per category one can scrape the internet freely. We model the defender cost as $c=n*(pl+cc)$ and the attacker cost as $c_{a}=n_{a}*(pl_{a}+cc_{a})$, where $n$ is the number of points to annotate, $pl$ is the per-label cost and $cc$ is the data collection cost. We assume that the attacker wants to succeed in the attack and thus attacker’s cost is less than that of a defender $n*(pl+cc)>n_{a}*(pl_{a}+cc_{a})$. First, we consider the attacker who extracts a model with DFME with 20 million queries. If we assume that the defender used Google Cloud to label their data, we find that the attacker’s labels have to cost less than $0.00012\$$:
$20M*pl_{a}<2,400\$\rightarrow pl_{a}<0.00012\$\;.$
The queries in DFME are synthetic, therefore there is no data collection cost ($cc_{a}=0$). Next, we consider a ME attack that only utilizes $5\%=3000$ datapoints of prior knowledge, and no additional queries of any sort. Here, attacker breaks even at $0.8\$$ per-label. Note that with 5% of data, as we show later in Section 5.1, it is indeed possible to extract a model, while remaining well within a reasonable budget in the current labeling market. At the same time, extraction with DFME stops being cost-effective when the attacker queries are even minimally priced by the defender.
This case study suggests that the claim over the cost-effectiveness of ME is not so trivial, and should be carefully considered in different use-cases. An attacker must estimate the cost elements defined above, namely sample complexity, data collection, and data labeling, in order to properly justify performing the attack. We now turn to the investigating the cost effectiveness of current ME attacks in the literature.
4 Methodology
@@
A common threat model assumes IND data is scarce and expensive to obtain, thus many attackers attempt to reduce their data collection costs through OOD queries. Note that this is the most common way for ME literature to reduce costs. But how might this work? Hyperbolizing, this implies that by asking questions about e.g. a ‘dog’, one may learn something about unrelated concept of a ‘building’. In this section, we first provide intuition for when such extraction is possible. We then build on this intuition and design an explicit mechanism to measure how much OOD queries can possibly contribute to attack performance.
@@
Warm-up: Linear classification Consider a linear classifier that splits the input space into two decision regions, and the corresponding decision boundary is defined by the line $y=\alpha x$. Here, $(x,y)$ denotes the two-dimensional input, and $\alpha$ is the single model parameter that determines the entire decision boundary, which we demonstrate in Figure 1. ME in this context is equivalent to estimating the decision boundary over the green IND region where $x\sim[20,80]$. The attacker is assumed to have the ability to query the model with arbitrary data but has no knowledge of where the IND region starts or ends. We assume that the depicted red region, $x\sim[0,20]\cup[80,100]$, is OOD. Note that it is only important to perform well on the task in the IND region, and neither the victim nor the attacker are concerned with making predictions for OOD inputs.
@@
The left-most plot represents the case of a naive victim model with the original decision boundary. Here ME reduces to estimating a single parameter $\alpha$, which can be achieved by using just 2 data points. Note that, in this case, the IND model can be extracted equally well using both IND or OOD data points, as the model behaviour is shared between both domains. In the middle case, we slightly increase the complexity of the decision boundary in the OOD area by introducing a new linear boundary with non-zero intercept in each region. Here, the OOD queries reveal less information over the IND behaviour, thus requiring the adversary to sample from both regions in order to faithfully learn the boundary. As the attacker does not know where the IND and OOD regions are, it must approximate a piece-wise linear boundary, i.e. the slopes and the intercepts of the new lines, in addition to the slope of the original IND boundary. This increases the required number of parameters to 5 and the minimal required sample size to 6, and consequently, the cost of ME. In the right-most case, we further increase the complexity by introducing two additional linear decision boundaries per region, resulting in a total of 9 parameters to be estimated using at least 10 data points.
@@
The above example demonstrates that while the IND area was not modified, adding complex and task-independent boundaries to OOD regions can significantly increase the attacker’s required capacity and sample complexity, thus making ME more costly. This behaviour is related to the locally independent nature of large models when performing regression or classification tasks. As sampling the boundary at one point does not necessarily reveal any information on the boundary IND, the attacker must explore the entire input domain. The main assumption here is that the attacker cannot know which areas are important. In other words, we find that ME attacks implicitly assume that the adversary has prior knowledge of the distribution to be able to reproduce the victim model’s predictions on the task of interest. Without such knowledge, the adversary is unable to tell whether making a query to the victim model will aid its goal of learning the true decision boundaries. Hence, we demonstrate that ME adversaries either require prior knowledge about the underlying distribution or need to submit a large number of queries to the victim model.
4.1 Sampling complexity intuition
4.2 On hardness of OOD detection
Recently, Tramer (2021) demonstrated that robust adversarial example detection is as hard as robust classification. Here, robustness of the detector refers to consistent detection over the $\epsilon$-radius around a data point, representing the maximal distance for an adversarial example. Fundamentally, Tramer demonstrates that robust detection over an $\epsilon$ region around a given (not necessarily adversarial) data point implies, albeit inefficient, an ability to successfully classify for at least $\epsilon/2$ region. One way to reason about the hardness of sampling from IND is to think about how informative the predictions of the victim model are on OOD queries. Simply put into Bayesian interpretation, Tramer says that an ability to compute $\mathbb{P}_{\mathcal{D}}(x=x_{i},y=y_{i})$, i.e. perform OOD detection robustly with knowledge of the labels, implies an ability to classify robustly $\mathbb{P}_{\mathcal{D}}(y=y_{i}|x=x_{i})$. Intuitively, that follows from the Bayes rule with $\mathbb{P}_{\mathcal{D}}(x=x_{i},y=y_{i})=\mathbb{P}_{\mathcal{D}}(y=y_{i}|x=x_{i})*\mathbb{P}_{\mathcal{D}}(x=x_{i})$, where $\mathbb{P}_{\mathcal{D}}(x=x_{i})$ represents prior knowledge over the true data distribution $\mathcal{D}$, i.e. the likelihood of $x_{i}$ coming from $\mathcal{D}$. In this paper we do not make assumptions about the labels – one can certainly imagine an attacker who has access to some labeled data. Do note however that label information is not necessary for a classic OOD detection i.e. checking if a given point comes from a given distribution $\mathbb{P}_{\mathcal{D}}(x=x_{i})$. There are two possibilities. First, if the attacker knows the labels of the points being queried, then the setting is exactly the same as the one considered by Tramer i.e. robust OOD detection is reducible to robust classification. Second, if the attacker is assumed to be capable of estimating $\mathbb{P}_{\mathcal{D}}(x=x_{i})$ robustly, reduction to classification without labels is impossible – the attacker would simply need to use the model as a labeling oracle.
Practically, this implies that with the OOD informativeness controlled, ME might not be the most efficient way for the attacker to achieve their objectives, given the capabilities they have access to. If they have the labels and an OOD detector, then they can already perform classification. If they do not have the labels, they can, at most, get the benefit of labeling their dataset, after which they get classification. If they are unable to solve the OOD detection problem, they would have to first build an OOD detector (which means collecting costly datapoints) since otherwise, OOD queries provide them no information about the IND region. In other words, once we take into account the OOD informativeness control, a model extraction attacker is only as good as their knowledge of the underlying dataset; but if this knowledge includes labels, often there is no need to extract models in the first place. This allows us to evaluate the performance of the attacker policy relative to the IND.
4.3 Out-of-Distribution instrumentation
In Section 5 we empirically evaluate our hypothesis that ME attacks make an implicit assumption about the usefulness of OOD queries. We show that when this assumption does not hold, and the OOD region is not indicative of the IND behaviour, ME attack success is reduced to being a function of prior knowledge. Due to space limitations, we only briefly describe the details of the model instrumentation, and refer the reader to Appendix F for full details.
Given the original victim model $\mathcal{V}_{o}$, we create a hybrid victim model $\mathcal{V}_{h}$ by combining $\mathcal{V}_{o}$ with an additional module $\mathcal{V}_{f}$ with different, or additional, decision boundaries. This additional model is used to provide predictions for OOD queries that differ from the predictions they would have gotten from the original model. For each query $x$, the hybrid model $\mathcal{V}_{h}$ applies some decision rule $R$ to classify $x$ as IND or OOD, and uses the corresponding model for prediction. We design the additional model $\mathcal{V}_{f}$ such that the decision boundaries of both models would have similar smoothness properties; thus, learning the “fake” boundaries is expected to be of the same level of difficulty, with nearly statistically indistinguishable output distributions. We discuss this further in Section F.4. For the decision rule $R$, we apply some pre-defined threshold $\tau$ over the prediction confidence of $\mathcal{V}_{o}$, with a softmax temperature of $2$ to better calibrate $R$. If the confidence is higher than $\tau$, the hybrid model returns $\mathcal{V}_{o}(x)$, otherwise it returns $\mathcal{V}_{f}(x)$. Prediction confidence serves as a naive OOD detector (Hendrycks and Gimpel, 2016; DeVries and Taylor, 2018) and represents one of the simplest settings for the attacker. More advanced OOD detectors will strictly improve the separation of IND and OOD queries, therefore decreasing the number of false negative queries, i.e. OOD queries that are classified IND and are provided with a meaningful prediction by $\mathcal{V}_{o}$. For the fake model $\mathcal{V}_{f}$ we fit a Gaussian Mixture Model (GMM) for each class, and split the input domain into centroids defined by an anchor point. These are used to “assign” queries to one of the GMMs for prediction with a random yet consistent class prediction around anchor point.
For a given query sample $x$, we compute its feature representation and find the nearest $L_{2}$ anchor point. We sample a fake prediction $\tilde{y}$ using the GMM corresponding to the chosen anchor point and return it as the fake model prediction $\mathcal{V}_{f}(x)=\tilde{y}$. The construction of the GMMs and their anchor points, discussed in detail in Appendix F, results in OOD samples being “wrongly” predicted while still exhibiting similar smoothness properties. Put altogether, we have presented here a method to control the usefulness of the victim model’s predictions on OOD queries – which will allow us to ablate its effect on the performance of ME.
We note that the principles of our instrumentation are similar to those used by the defense proposed by Kariyappa and Qureshi (2020), however the exact construction of $\mathcal{V}_{f}$ differs, where Kariyappa and Qureshi (2020) train a separate model to provide incorrect responses, while we sample statistically plausible ones.
5 Evaluation
In the next subsections sections we empirically answer the following five questions:
  • (Section 5.1) Does model extraction work? Yes, model extraction definitively works in that it is possible to approximate a blackbox model from just queries, although not all queries are equally useful.
  • (Section 5.2) Is it possible to extract a model using small number of queries? Sometimes, assuming the attacker has access to some prior knowledge over the distribution and can choose a limited number of high-quality queries that provide task-informative responses.
  • (Section 5.3) Can model extraction be used to reduce the data costs? Sometimes, assuming that OOD responses provide task-informative answers, data costs can be reduced.
  • (Section 5.4) Can model extraction be used both to reduce the data costs and at the same time use small number of queries? No, if OOD only reveals limited information about the IND behaviour, a ME attacker must either have prior knowledge of the distribution and the ability to sample IND samples or have a very large query budget and traverse the input domain.
  • (Section 5.5) Can model extraction be used to reduce data labeling costs? Sometimes; we find that in practice the victim model provides no additional benefits other than serving as a labeling oracle. For some use cases, there are no more cost-effective sources for obtaining labels; for example, for medical data.
Experimental Setting.